OpenStudy (bigbosssaint21):

Solve 2x2 -10x + 3 = 0

OpenStudy (anonymous):

Isolate x.

OpenStudy (anonymous):

2x^2 - 10x + 3 = 0 2x^2 - 10x = -3

OpenStudy (anonymous):

2x^2 = 10x - 3

OpenStudy (anonymous):

Wait, messed up. derp

OpenStudy (anonymous):

The answer is 4.6794494... Just round it to 4.7

OpenStudy (anonymous):

Sorry x approximately equals 4.7

OpenStudy (anonymous):

@tatiana__celeste__ you should read the code of conduct, we at OpenStudy do not just give out answers. This is a tutoring website, you basically just let yourself be a calculator.

OpenStudy (jdoe0001):

@Bigbosssaint21 have you covered the quadratic formula yet?